OpenStudy (daisy.xoxo619):

A den is 8 feet longer that it is wide. if the dens area is 308 square feet what are the dimensions of the room

OpenStudy (mathmale):

this being algebra, you should choose letters of the alphabet to represent the width and then the length of the den. Let the width be ... ? Let the length be ... ? show how the length is related to the width.

OpenStudy (daisy.xoxo619):

how can i set up the equation?

OpenStudy (mathmale):

Representation of your unknowns comes before setting up any equation. Choose a letter, now, to represent the width of the den.

OpenStudy (daisy.xoxo619):

w for width l for length

OpenStudy (mathmale):

Let x = width of den. Since the length is 8 feet longer than the width, the length is represented by x + 8. Given the L and the W of any rectangle, how would you find the area? Write out the area formula, please.

OpenStudy (daisy.xoxo619):

A=WL A=X*X+8 Is that right?

OpenStudy (mathmale):

Yes, but you do have to enclose that "x+8" inside parentheses. Do that now, please. A =

OpenStudy (daisy.xoxo619):

A=x+(x+8)

OpenStudy (mathmale):

But you multiplied in your first try, but now are adding. Area is L * W, not L + W. Try writing the equation again, please. A = 308 sq ft =

OpenStudy (daisy.xoxo619):

A=308 sq ft= x*(x+8)

OpenStudy (mathmale):

good. Please multiply out the right hand side of this equation.

OpenStudy (daisy.xoxo619):

x^2+8x

OpenStudy (mathmale):

x^2+8x=308. Good. How would y ou solve this quadratic equation?

OpenStudy (daisy.xoxo619):

subtract 308 from both sides

OpenStudy (mathmale):

yes

OpenStudy (daisy.xoxo619):

I'm going to use quadratic formula, hold on please

OpenStudy (daisy.xoxo619):

does x=14, -22?

OpenStudy (mathmale):

-22 could not be correct, because length is never negative. That leaves you with 14. Check your work. x(x+8) must equal 308. Substitute 14 for x. Is the resulting equatin true or false?

OpenStudy (daisy.xoxo619):

ok one second

OpenStudy (daisy.xoxo619):

yes i got 308

OpenStudy (mathmale):

Very good. Great. That means your work is correct and you've solved this algebra problem. Keep in mind the importance of being able to represent unknown quantities with letters of the alphabet, as we did with W = x and L = x + 8. Over and out, good night!

OpenStudy (daisy.xoxo619):

thank you so much

OpenStudy (mathmale):

You're welcome, Daisy! Bye!